# Calibrating LLM Confidence by Probing Perturbed Representation Stability

### v1

@@ Abstract
Reza Khanmohammadi1, Erfan Miahi2, Mehrsa Mardikoraem2, Simerjot Kaur3, Ivan Brugere3, Charese H. Smiley3, Kundan Thind4, Mohammad M. Ghassemi1 1Michigan State University  2Independent AI Researcher 3JPMorgan AI Research  4Henry Ford Health Correspondence: khanreza@msu.edu
@@ 1 Introduction
Despite their impressive performance, large language models (LLMs) consistently struggle with confidence calibration Guo et al. (2017); Geng et al. (2024). Their confidence—the model’s internally estimated probability that a given response is correct—frequently misaligns with actual outcomes: LLMs often assign high confidence to wrong answers and low confidence to right ones. This unreliability is particularly acute in high-stakes domains like medicine, finance, and law. For example, in a critical medical task like symptom extraction for cancer toxicity assessment, even if an LLM often produces correct information, it might do so with inappropriately low confidence, or conversely, express high confidence for incorrect outputs. If such confidence scores are not dependable guides to actual correctness, human experts may be forced to meticulously review every LLM-generated instance, significantly diminishing the practical benefits of automation. Accurate confidence estimation for each specific response is therefore essential, as it provides a vital mechanism for managing risk, enabling users to prioritize human oversight, selectively rely on LLM outputs, and ultimately foster more responsible and effective LLM integration.
Existing approaches to LLM confidence estimation include direct self-evaluation Kadavath et al. (2022), post-hoc adjustments Jiang et al. (2021), internal state probing with lightweight classifiers Azaria and Mitchell (2023); Liu et al. (2024), and model fine-tuning Kapoor et al. (2024b). These methods often struggle to consistently deliver on multiple desirable properties simultaneously, namely achieving strong calibration (e.g., low Expected Calibration Error (ECE)) and high discriminative power (e.g., high Area Under the Precision-Recall Curve (AUCPR) or Area Under the Receiver Operating Characteristic Curve (AUROC)) while maintaining computational efficiency and robust generalizability across the diverse set of LLM architectures and families. Many methods excel in some of these desirable properties but make trade-offs in others; for instance, fine-tuning approaches like Calibration-Tuning (CT) Kapoor et al. (2024b) often achieve strong calibration in ECE but may not consistently lead in discriminative metrics like AUROC, while lightweight methods such as LitCab Liu et al. (2024) can demonstrate strong AUROC but sometimes show variable ECE performance across different LLM families. This leaves a need for more holistically effective solutions.
@@ 3.1 Probing Internal Stability
Adversarial Perturbation Trajectory Direction  To define a systematic perturbation trajectory that challenges the LLM’s generation of the observed token $t_{i}$, we utilize the gradient of the loss associated with $t_{i}$ with respect to its generating hidden state $H_{0}^{(i)}$. Let $P(t_{i}|H_{0}^{(i)})$ be the probability of the token $t_{i}$ given $H_{0}^{(i)}$. We define the loss as the negative log-likelihood: $\mathcal{L}^{(i)}=-\log P(t_{i}|H_{0}^{(i)})$. The Jacobian vector $J^{(i)}=\nabla_{H_{0}^{(i)}}\mathcal{L}^{(i)}$ then indicates the direction in the hidden state space where this loss $\mathcal{L}^{(i)}$ increases most rapidly; equivalently, this is the direction where the probability of token $t_{i}$ decreases most steeply. We normalize this vector to obtain the unit direction $d^{(i)}=J^{(i)}/||J^{(i)}||_{2}$. If $J^{(i)}$ is a zero vector, $d^{(i)}$ is also set to zero. Perturbing along this direction $d^{(i)}$ is an adversarial act [^1] Our use of “adversarial” here refers to targeted, gradient-informed perturbations designed to systematically probe representational stability by challenging the generation of token $t_{i}$. This is distinct from adversarial attacks aimed at finding minimal input perturbations to cause misclassification or from adversarial training regimens. aimed at making the original token $t_{i}$ less likely. This contrasts with standard LLM training where one steps in the negative gradient direction (e.g., $-\nabla\mathcal{L}$) to reduce loss for a target token. Here, by moving along the positive gradient of $\mathcal{L}^{(i)}$, we are adversarially probing the stability of the LLM’s initial choice $t_{i}$ by actively trying to dislodge it.
Iterative Adversarial Perturbation  We then explore the stability of $H_{0}^{(i)}$ by applying $S$ discrete adversarial perturbations along the direction $d^{(i)}$. The maximum extent of this exploration is defined by a radius $\epsilon_{\text{max}}$ (in our experiments, $\epsilon_{\text{max}}=20.0$ and $S=5$). The $s$-th perturbation magnitude is $\epsilon_{s}=s\cdot(\epsilon_{\text{max}}/S)$, for $s\in\{1,\ldots,S\}$. The $s$-th perturbed hidden state is:
@@ 3.3 Confidence Classification Architecture
Both $E_{\text{MC}}$ and $E_{\text{OE}}$ are pre-trained using a Max-Margin contrastive loss. This choice of loss is aimed at learning discriminative embeddings, a strategy also found effective in other confidence estimation works such as Liu et al. (2024), though our feature generation process and overall architecture are distinct. The objective of this pre-training is to map features from correctly answered questions to regions in the embedding space that are separable from those associated with incorrect answers, supervised by the ground truth correctness of $A$.
@@
Training Procedure  Following the contrastive pre-training of the projection network, the projection network ($E_{\text{MC}}$ or $E_{\text{OE}}$) and the classification head $C$ are jointly fine-tuned. This stage employs a standard cross-entropy loss, again supervised by the ground truth correctness of answer $A$.
@@ 4 Experimental Setup
We empirically evaluate the effectiveness of our proposed confidence estimation method.
Languge Models  Our experiments utilize a range of contemporary decoder-only LLMs, with parameter sizes from 8B to 32B, encompassing three distinct model families to assess broader applicability. Specifically, we include: Meta-Llama-3.1-8B-Instruct Grattafiori et al. (2024) (governed by the Llama 3.1 Community License Agreement), Qwen2.5-14B-Instruct Qwen et al. (2025) (Apache License 2.0), Mistral-Small-24B-Instruct-2501 (Apache License 2.0), and Qwen2.5-32B-Instruct Qwen et al. (2025) (Apache License 2.0). We use the unsloth Daniel Han and team (2023) versions of these models to leverage their efficient, non-quantized model handling capabilities. All experiments, including hidden state extraction and model training, were conducted using bfloat16 precision to ensure manageability across the diverse models and computational tasks. Our implementations rely on the HuggingFace Transformers Wolf et al. (2020) and PyTorch Paszke et al. (2019) libraries.
@@
Datasets  For training and validating our confidence estimation models, we utilize the CT-CHOICE and CT-OE datasets for multiple-choice and open-ended question-answering formats, respectively. These datasets, generated following the exact methodology detailed by Kapoor et al. (2024b) (Apache License 2.0), comprise a large collection of commonly used question-answering datasets from the literature. To assess generalization and performance, we evaluate on tasks from the Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2021) (MIT License). We created multiple-choice and open-ended versions of these tasks, namely MMLU-CHOICE and MMLU-OE, using the same data processing approach as Kapoor et al. (2024b) to ensure consistency. Additionally, we employ MMLU-PRO-CHOICE (Apache License 2.0), a multiple-choice version of the MMLU-Pro dataset Wang et al. (2024), for further rigorous testing. All dataset instances across training, validation, and testing incorporate 5-shot exemplars within the input prompt $P$ to contextualize the LLMs. Additional details on the dataset characteristics, response generation process, and labeling procedures are provided in Appendix C.
@@
Training Details  To ensure fair comparisons, training configurations were kept consistent across all methods, including baselines. The main classification/fine-tuning stage for all models involved a total of 10,000 training steps. For our proposed method, the contrastive feature projection network ($E_{\text{MC}}$ or $E_{\text{OE}}$) was pre-trained for 5,000 steps. Subsequently, the confidence classification model was trained for an additional 5,000 steps. Key hyperparameters for the AdamW optimizer Loshchilov and Hutter (2019), such as a learning rate of $1\times 10^{-4}$, were aligned with those reported by Kapoor et al. (2024b). Training was conducted with a batch size of 32. A weight decay of 0.1 was uniformly applied across all training stages and methods.
@@
Baselines  We compare our method (CCPS) against a comprehensive set of established confidence estimation techniques. These include P(True) and P(IK) Kadavath et al. (2022), Logit Temperature Scaling (LTS) Jiang et al. (2021), Instruction Tuning (IT) Wei et al. (2022) on the uncertainty query, SAPLMA Azaria and Mitchell (2023) (with variants SAPLMA-F, SAPLMA-M, and SAPLMA-UM corresponding to different layer inputs), Calibration Tuning (CT) Kapoor et al. (2024b), and LitCab Liu et al. (2024). Detailed descriptions of these baseline methods are provided in Appendix D, and the architectural specifics of our CCPS model, including details of the hyperparameter search process that determined these architectures, are detailed in Appendix E. Information regarding the computational setup and resources utilized for all methods is available in Appendix F. Furthermore, a comparative analysis of the additional trainable parameters introduced by each method is presented in Appendix G, underscoring the parameter efficiency of our CCPS approach.
Evaluation Metrics  We evaluate our confidence estimation method using established metrics. For assessing calibration, we employ the ECE and the Brier Score. We also measure our confidence scores’ discriminative performance using ACC, AUCPR, and AUROC. Detailed definitions and formulations for these metrics are provided in Appendix H.
@@ 5 Results
On the standard multiple-choice benchmark, MMLU-CHOICE, CCPS consistently achieves superior performance across all four base LLMs. For instance, ECE scores for CCPS are typically in
Method ECE $\downarrow$ BRIER $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$
@@
Meta-Llama-3.1-8B-Instruct
LitCab 10.9 18.1 73.2 84.0 77.1
CT 10.7 21.1 67.8 74.2 62.8
CCPS 6.5 17.1 73.4 84.1 77.1
Qwen2.5-14B-Instruct
LitCab 45.6 20.0 78.3 83.7 65.3
CT 12.1 17.0 78.6 84.7 64.8
CCPS 6.3 13.1 80.2 92.1 81.6
Mistral-Small-24B-Instruct-2501
LitCab 13.5 15.1 79.5 91.5 78.2
CT 8.2 15.5 79.6 83.3 56.5
CCPS 5.8 11.5 83.0 93.1 83.3
Qwen2.5-32B-Instruct
LitCab 43.2 15.9 82.6 87.9 67.2
CT 45.2 46.9 37.2 84.3 51.6
CCPS 6.3 10.8 84.1 94.1 82.8
@@
Meta-Llama-3.1-8B-Instruct
LitCab 16.6 24.7 66.1 51.7 63.6
CT 21.5 29.8 50.4 43.7 57.3
CCPS 4.5 20.0 70.4 55.2 67.9
Qwen2.5-14B-Instruct
LitCab 49.7 38.3 55.3 66.2 68.0
CT 20.4 28.7 55.6 59.4 56.6
CCPS 4.2 20.1 69.2 75.8 74.0
Mistral-Small-24B-Instruct-2501
LitCab 32.2 34.6 57.0 66.2 60.1
CT 17.8 27.4 58.2 60.1 54.3
CCPS 4.5 18.6 71.3 79.5 77.2
Qwen2.5-32B-Instruct
LitCab 48.4 33.7 60.8 72.7 70.3
CT 38.0 41.6 44.8 60.5 49.9
CCPS 4.6 18.5 71.8 82.4 77.8
@@
Meta-Llama-3.1-8B-Instruct
LitCab 8.8 22.5 65.3 46.2 66.0
CT 8.8 21.1 65.3 48.9 70.9
CCPS 8.0 20.2 69.5 49.4 69.3
Qwen2.5-14B-Instruct
LitCab 34.4 37.0 49.4 56.8 62.5
CT 9.4 22.6 63.4 61.7 69.3
CCPS 6.7 22.5 63.6 59.0 66.6
Mistral-Small-24B-Instruct-2501
LitCab 11.2 24.6 60.2 60.5 66.4
CT 10.8 22.8 62.2 60.7 68.2
CCPS 6.8 20.8 67.6 64.7 71.4
Qwen2.5-32B-Instruct
LitCab 28.4 33.2 52.7 60.2 62.3
CT 22.9 31.1 57.1 52.9 56.3
CCPS 8.7 23.3 62.6 62.0 66.4
Table 1: Average performance on MMLU variants across tasks per LLM. Arrows indicate whether lower ($\downarrow$) or higher ($\uparrow$) values are better. All values are percentages. Best values per method-block are bolded.
the range of 5.8-6.5%, representing substantial reductions compared to both LitCab and CT, which often exhibit much higher ECEs (e.g., LitCab’s ECE of 45.6% and CT’s 45.2% on Qwen2.5-14B and Qwen2.5-32B respectively, against CCPS’s 6.3% on both). CCPS shows similar gains in Brier score and discriminative metrics like AUCPR and AUROC, often matching or outperforming baselines.
@@
In the open-ended generation setting (MMLU-OE), CCPS generally maintains strong calibration, consistently achieving the best ECE and Brier scores, especially with larger models like Mistral-24B and Qwen2.5-32B where it leads across all metrics. For smaller models on MMLU-OE, while CCPS leads in calibration, CT demonstrates competitive discriminative performance in AUCPR and AUROC (e.g., for Llama-3.1-8B, CT’s AUROC is 70.9% vs. CCPS’s 69.3%; for Qwen2.5-14B, CT leads in AUCPR and AUROC). However, CCPS’s calibration advantage remains evident, for example, achieving an ECE of 6.7% with Qwen2.5-14B compared to CT’s 9.4%.
In summary, CCPS consistently delivers substantial improvements in confidence estimation, excelling in both calibration and the ability to discriminate between correct and incorrect responses across diverse LLMs and task formats, particularly on challenging multiple-choice benchmarks. The findings in Table 1 are further detailed in Appendix I, which includes comprehensive results for all baselines (mean and standard deviation scores, comparative bar charts, per-task breakdowns, and feature importance analyses).

### v2

@@ Abstract
Reza Khanmohammadi1thanks: Corresponding author: khanreza@msu.edu, Erfan Miahi2, Mehrsa Mardikoraem2, Simerjot Kaur3, Ivan Brugere3, Charese H. Smiley3, Kundan Thind4, Mohammad M. Ghassemi1 1Michigan State University  2Independent AI Researcher 3JPMorgan AI Research  4Henry Ford Health 1 {khanreza, ghassem3}@msu.edu2 {mhi.erfan1, mardikorm}@gmail.com 3 {simerjot.kaur, ivan.brugere, charese.h.smiley}@jpmchase.com4 kthind1@hfhs.org
@@ 1 Introduction
Despite their impressive performance, large language models (LLMs) consistently struggle with confidence calibration Guo et al. (2017); Geng et al. (2024). Their confidence—the model’s internally estimated probability that a given response is correct—frequently misaligns with actual outcomes: LLMs often assign high confidence to wrong answers and low confidence to right ones. This unreliability is particularly acute in high-stakes domains like medicine, finance, and law. For example, in a medical task like symptom extraction for cancer toxicity assessment, even if an LLM often produces correct information, it might do so with inappropriately low confidence, or conversely, express high confidence for incorrect outputs. If such confidence scores are not dependable guides to actual correctness, human experts may be forced to meticulously review every LLM-generated instance, significantly diminishing the practical benefits of automation. Accurate confidence estimation for each specific response is therefore essential, as it provides a vital mechanism for managing risk, enabling users to prioritize human oversight, selectively rely on LLM outputs, and foster more responsible and effective LLM integration.
Existing approaches to LLM confidence estimation include direct self-evaluation Kadavath et al. (2022), post-hoc adjustments Jiang et al. (2021), internal state probing with lightweight classifiers Azaria and Mitchell (2023); Liu et al. (2024), and model fine-tuning Kapoor et al. (2024b). These methods often struggle to consistently deliver on multiple desirable properties simultaneously, namely achieving strong calibration (e.g., low Expected Calibration Error (ECE)) and high discriminative power (e.g., high Area Under the Precision-Recall Curve (AUCPR) or Area Under the Receiver Operating Characteristic Curve (AUROC)) while maintaining computational efficiency and generalizability across the diverse set of LLM architectures and families. Many methods excel in some of these desirable properties but make trade-offs in others; for instance, fine-tuning approaches like Calibration-Tuning (CT) Kapoor et al. (2024b) often achieve strong calibration in ECE but may not consistently lead in discriminative metrics like AUROC, while lightweight methods such as LitCab Liu et al. (2024) can demonstrate strong AUROC but sometimes show variable ECE performance across different LLM families. This leaves a need for more holistically effective solutions.
@@ 3 Method
In this work, we use the term adversarial in a restricted sense: it refers to targeted, gradient-informed perturbations that are designed to systematically probe representational stability by challenging the generation of token $t_{i}$. This usage is distinct from adversarial attacks aimed at inducing misclassification with minimal input changes, or from adversarial training schemes intended to improve model robustness.
@@ 3.1 Probing Internal Stability
[…]
Iterative Adversarial Perturbation  We then explore the stability of $H_{0}^{(i)}$ by applying $S$ discrete adversarial perturbations along the direction $d^{(i)}$. The maximum extent of this exploration is defined by a radius $\epsilon_{\text{max}}$ and the number of steps is $S$. We deliberately used a fixed set of hyperparameters ($\epsilon_{\text{max}}=20.0,S=5$) across all models and benchmarks to test the method’s robustness. These values were chosen based on empirical observation; $\epsilon_{\text{max}}=20.0$ was a practical upper bound where the integrity of the LLM’s output distribution began to degrade into noise or "gibberish" consistently across all models, while $S=5$ was selected as a pragmatic trade-off between capturing the trajectory’s dynamics and managing the computational cost. The $s$-th perturbation magnitude is $\epsilon_{s}=s\cdot(\epsilon_{\text{max}}/S)$, for $s\in\{1,\ldots,S\}$. The $s$-th perturbed hidden state is:
@@ 3.3 Confidence Classification Architecture
Both $E_{\text{MC}}$ and $E_{\text{OE}}$ are pre-trained using a Max-Margin contrastive loss. Specifically, given a correct answer embedding $e^{+}$ and an incorrect answer embedding $e^{-}$ for the same question, the loss encourages the distance between $e^{+}$ and $e^{-}$ to exceed a margin $\gamma>0$, formulated as
$\mathcal{L}_{\text{max-margin}}=\max\!\big(0,\,\gamma-(\|e^{+}-e^{-}\|_{2}-\|e^{+}-e^{+}\|_{2})\big)$
This objective pushes embeddings of correct answers closer together while enforcing separation from incorrect ones. The choice of loss is aimed at learning discriminative embeddings, a strategy also found effective in other confidence estimation works such as Liu et al. (2024). The objective of this pre-training is to map features from correctly answered questions to regions in the embedding space that are separable from those associated with incorrect answers, supervised by the ground truth correctness of $A$.
@@
Training Procedure  Following the contrastive pre-training of the projection network, the projection network ($E_{\text{MC}}$ or $E_{\text{OE}}$) and the classification head $C$ are jointly fine-tuned. This stage employs a standard cross-entropy loss, again supervised by the ground truth correctness of answer $A$. Further implementation details of both $E_{\text{MC}}$ and $E_{\text{OE}}$, including the results of our hyperparameter search and the finalized layer configurations, are provided in Appendix E.
@@ 4 Experimental Setup
This section details the experimental setup designed to empirically evaluate CCPS. Our evaluation framework provides a comprehensive and rigorous comparison, benchmarking CCPS against a wide array of recent confidence estimation methods. This benchmark is conducted across four modern LLMs of varying architectures and scales, on knowledge-based question-answering (QA) datasets including MMLU and MMLU-Pro in both multiple-choice and open-ended formats. The following subsections detail the specific language models, datasets, training configurations, baselines, and evaluation metrics employed in our study.
@@
Datasets  For training and validating our confidence estimation models, we utilize the CT-CHOICE and CT-OE datasets for multiple-choice and open-ended QA formats, respectively. These datasets, generated following the exact methodology detailed by Kapoor et al. (2024b) (Apache License 2.0), comprise a large collection of commonly used QA datasets from the literature. To assess generalization and performance, we evaluate on tasks from the Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2021) (MIT License). We created multiple-choice and open-ended versions of these tasks, namely MMLU-CHOICE and MMLU-OE, using the same data processing approach as Kapoor et al. (2024b) to ensure consistency. Additionally, we employ MMLU-PRO-CHOICE (Apache License 2.0), a multiple-choice version of the MMLU-Pro dataset Wang et al. (2024), for further rigorous testing. All dataset instances across training, validation, and testing incorporate 5-shot exemplars within the input prompt $P$ to contextualize the LLMs. Additional details on the dataset characteristics, response generation process, and labeling procedures are provided in Appendix C.
@@
Training Details  The full architectural details of our projection networks and classification head are provided in Appendix E. To ensure fair comparisons, training configurations were kept consistent across all methods, including baselines. The main classification/fine-tuning stage for all models involved a total of 10,000 training steps. For our proposed method, the contrastive feature projection network ($E_{\text{MC}}$ or $E_{\text{OE}}$) was pre-trained for 5,000 steps. Subsequently, the confidence classification model was trained for an additional 5,000 steps. Key hyperparameters for the AdamW optimizer Loshchilov and Hutter (2019), such as a learning rate of $1\times 10^{-4}$, were aligned with those reported by Kapoor et al. (2024b). Training was conducted with a batch size of 32. A weight decay of 0.1 was uniformly applied across all training stages and methods.
@@
Baselines  We compare our method (CCPS) against a comprehensive set of established confidence estimation techniques. These include P(True) and P(IK) Kadavath et al. (2022), Logit Temperature Scaling (LTS) Jiang et al. (2021), Instruction Tuning (IT) Wei et al. (2022) on the uncertainty query, SAPLMA Azaria and Mitchell (2023) (with variants SAPLMA-F, SAPLMA-M, and SAPLMA-UM corresponding to different layer inputs), Calibration Tuning (CT) Kapoor et al. (2024b), and LitCab Liu et al. (2024). Detailed descriptions of these baseline methods are provided in Appendix D. Information regarding the computational setup and resources utilized for all methods is available in Appendix F. Furthermore, a comparative analysis of the additional trainable parameters introduced by each method is presented in Appendix G, underscoring the parameter efficiency of our CCPS approach.
Evaluation Metrics  We focus on two primary metrics in the main text. For calibration, we use the Expected Calibration Error (ECE), which measures how well predicted confidences align with actual accuracies. Intuitively, if a model assigns 70% confidence to a set of answers, then about 70% of those answers should be correct. To compute this, the $n$ samples are partitioned into $b=10$ equally spaced bins $\{B_{j}\}_{j=1}^{b}$, and we compare the average predicted confidence $\text{conf}(B_{j})$ against the empirical accuracy $\text{acc}(B_{j})$ within each bin:
$\text{ECE}=\sum_{j=1}^{b}\frac{|B_{j}|}{n}\,\big|\text{conf}(B_{j})-\text{acc}(B_{j})\big|.$
Smaller values indicate that predicted probabilities more faithfully reflect true correctness rates.
In addition, we report the Brier Score, which directly measures the squared error between the predicted confidence $p_{k}$ and the ground-truth outcome $o_{k}\in\{0,1\}$ for each sample $k$:
$\text{Brier}=\frac{1}{n}\sum_{k=1}^{n}(p_{k}-o_{k})^{2}.$
This metric captures both calibration and the sharpness of predictions, with lower scores reflecting more reliable and informative confidence estimates.
Additional classification-oriented metrics—including ACC, AUCPR, and AUROC—are reported and formally defined in Appendix H.
@@ 5 Results
On the standard multiple-choice benchmark, MMLU-CHOICE, CCPS consistently achieves superior performance across all four base LLMs. For instance, ECE scores for CCPS are typically in the range of 5.8-6.5%, representing substantial reductions compared to both LitCab and CT, which often exhibit much higher ECEs (e.g., LitCab’s ECE of 45.6% and CT’s 45.2% on Qwen2.5-14B and Qwen2.5-32B respectively, against CCPS’s 6.3% on both). CCPS shows similar gains in Brier score and discriminative metrics like AUCPR and AUROC, often matching or outperforming baselines.
@@
Meta-Llama-3.1-8B-Instruct Qwen2.5-14B-Instruct
Method ECE $\downarrow$ BRIER $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$ Method ECE $\downarrow$ BRIER $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$
LitCab 10.9 18.1 73.2 84.0 77.1 LitCab 45.6 20.0 78.3 83.7 65.3
CT 10.7 21.1 67.8 74.2 62.8 CT 12.1 17.0 78.6 84.7 64.8
CCPS 6.5 17.1 73.4 84.1 77.1 CCPS 6.3 13.1 80.2 92.1 81.6
Mistral-Small-24B-Instruct-2501 Qwen2.5-32B-Instruct
Method ECE $\downarrow$ BRIER $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$ Method ECE $\downarrow$ BRIER $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$
LitCab 13.5 15.1 79.5 91.5 78.2 LitCab 43.2 15.9 82.6 87.9 67.2
CT 8.2 15.5 79.6 83.3 56.5 CT 45.2 46.9 37.2 84.3 51.6
CCPS 5.8 11.5 83.0 93.1 83.3 CCPS 6.3 10.8 84.1 94.1 82.8
@@
Meta-Llama-3.1-8B-Instruct Qwen2.5-14B-Instruct
Method ECE $\downarrow$ BRIER $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$ Method ECE $\downarrow$ BRIER $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$
LitCab 16.6 24.7 66.1 51.7 63.6 LitCab 49.7 38.3 55.3 66.2 68.0
CT 21.5 29.8 50.4 43.7 57.3 CT 20.4 28.7 55.6 59.4 56.6
CCPS 4.5 20.0 70.4 55.2 67.9 CCPS 4.2 20.1 69.2 75.8 74.0
Mistral-Small-24B-Instruct-2501 Qwen2.5-32B-Instruct
Method ECE $\downarrow$ BRIER $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$ Method ECE $\downarrow$ BRIER $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$
LitCab 32.2 34.6 57.0 66.2 60.1 LitCab 48.4 33.7 60.8 72.7 70.3
CT 17.8 27.4 58.2 60.1 54.3 CT 38.0 41.6 44.8 60.5 49.9
CCPS 4.5 18.6 71.3 79.5 77.2 CCPS 4.6 18.5 71.8 82.4 77.8
@@
Meta-Llama-3.1-8B-Instruct Qwen2.5-14B-Instruct
Method ECE $\downarrow$ BRIER $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$ Method ECE $\downarrow$ BRIER $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$
LitCab 8.8 22.5 65.3 46.2 66.0 LitCab 34.4 37.0 49.4 56.8 62.5
CT 8.8 21.1 65.3 48.9 70.9 CT 9.4 22.6 63.4 61.7 69.3
CCPS 8.0 20.2 69.5 49.4 69.3 CCPS 6.7 22.5 63.6 59.0 66.6
Mistral-Small-24B-Instruct-2501 Qwen2.5-32B-Instruct
Method ECE $\downarrow$ BRIER $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$ Method ECE $\downarrow$ BRIER $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$
LitCab 11.2 24.6 60.2 60.5 66.4 LitCab 28.4 33.2 52.7 60.2 62.3
CT 10.8 22.8 62.2 60.7 68.2 CT 22.9 31.1 57.1 52.9 56.3
CCPS 6.8 20.8 67.6 64.7 71.4 CCPS 8.7 23.3 62.6 62.0 66.4
Table 1: Average performance on MMLU variants across tasks per LLM. Arrows indicate whether lower ($\downarrow$) or higher ($\uparrow$) values are better. All values are percentages. Best values per method-block are bolded. For brevity, only the two best-performing baselines are shown here; full results are provided in Appendix J.
@@
In the open-ended generation setting (MMLU-OE), CCPS generally maintains strong calibration, consistently achieving the best ECE and Brier scores, especially with larger models like Mistral-24B and Qwen2.5-32B where it leads across all metrics. For smaller models on MMLU-OE, while CCPS leads in calibration, CT demonstrates competitive discriminative performance in AUCPR and AUROC (e.g., for Llama-3.1-8B, CT’s AUROC is 70.9% vs. CCPS’s 69.3%; for Qwen2.5-14B, CT leads in AUCPR and AUROC). However, CCPS’s calibration advantage remains evident, for example, achieving an ECE of 6.7% with Qwen2.5-14B compared to CT’s 9.4%. To further assess the cross-domain robustness of CCPS, we conducted an additional evaluation on the specialized MedMCQA benchmark; these results are detailed in Appendix I.
In summary, CCPS consistently delivers substantial improvements in confidence estimation, excelling in both calibration and the ability to discriminate between correct and incorrect responses across diverse LLMs and task formats, particularly on challenging multiple-choice benchmarks. The findings in Table 1 are further detailed in Appendix J, which includes comprehensive results for all baselines (mean and standard deviation scores, comparative bar charts, per-task breakdowns, and feature importance analyses).